Multiple Choice: Justify Answer

3. If income is held constant and the wage rate increases, the desired hours of work will

a. decrease.

b. increase.

c. stay the same.

d. change ambiguously.
Answer #1

If income is held constant and the wage rate increases, the desired hours of work will DECREASE.

Let us understand this with an example.
Suppose an employer gives its employees a wage rate of $20 per hour. The income earnt for working 5 hours would, therefore, be $20*5hrs= $100 (20*5=100)
Now suppose this income of $100 per day is held constant. With this the wage rate increases to $40 per hour.

Since the employee cannot earn an income of more than $100 dollars because it is held constant, the working hours of the employee will decrease since the wage rate has increased. Let us calculate the number of hours a worker will work in a day to earn $100 with the new wage rate of $40 per hour.

Assume the new number of working hours= x

Therefore,

$40*x=$100
x=100/40
x=2.5

Therefore under the new wage rate of $40 dollars per hour the working hours will reduce from 5 hours (when it was $20 per hour) to 2.5 hours when the income of $100 is kept constant.

Hence, it is proved that with an increase in wage rate while keeping income constant, the desired number of working hours will Decrease.